# HumAIne-Chatbot: Real-Time Personalized Conversational AI via Reinforcement Learning

### v1

@@ Abstract
Current conversational AI systems often provide generic, one-size-fits-all interactions that overlook individual user characteristics and lack adaptive dialogue management. To address this gap, we introduce HumAIne-chatbot, an AI-driven conversational agent that personalizes responses through a novel user profiling framework. The system is pre-trained on a diverse set of GPT-generated virtual personas to establish a broad prior over user types. During live interactions, an online reinforcement learning agent refines per-user models by combining implicit signals (e.g., typing speed, sentiment, engagement duration) with explicit feedback (e.g., likes and dislikes). This profile dynamically informs the chatbot’s dialogue policy, enabling real-time adaptation of both content and style. To evaluate the system, we conducted controlled experiments with 50 synthetic personas across multiple conversation domains. Results showed consistent improvements in user satisfaction, personalization accuracy, and task achievement when personalization features were enabled. Statistical analysis confirmed significant differences between personalized and non-personalized conditions, with large effect sizes across key metrics. These findings highlight the effectiveness of AI-driven user profiling and provide a strong foundation for future real-world validation.
@@ III-C4 AI-Driven User Profiler with Reference Data Integration
Phase I: Virtual Persona Pre-training
We leverage the Anthology of GPT-generated backstories to simulate a diverse set of virtual personas. Each backstory encodes rich demographic and behavioral traits (age, occupation, language style), conditioning an LLM to "play" a consistent user archetype [32]. By engaging these personas in multi-turn dialogues, we collect a large synthetic corpus with "ground-truth" profile labels. The profiler is first trained in a supervised manner on this corpus, learning to infer latent user attributes and preferences from conversational cues.
This approach aligns with the VirtualXAI framework, which uses GPT-generated personas for qualitative assessment of explainability methods [33].
@@
We adopt Proximal Policy Optimization (PPO) [36] due to its stability and sample efficiency in high-dimensional, continuous action spaces. PPO updates policy parameters via clipped objective functions, ensuring gradual behavioral shifts—essential for maintaining user trust during personalization. The actor-critic architecture computes advantage estimates using a value network baseline, reducing variance in gradient estimates.
@@
Our two-stage framework builds on recent work combining persona simulation and RL. Wang et al. [35] demonstrate efficacy of a supervised-to-RL pipeline for realistic user simulators, while Li et al. [34] show that policy gradient methods improve dialogue coherence and engagement.
@@ IV Evaluation Strategy
To validate HumAIne-chatbot, we adopted a virtual persona–based evaluation approach inspired by VirtualXAI framework [37] that has used the approached proposed by [13]. Instead of human recruitment, we generated 50 diverse virtual personas with systematically varied demographics (18–65+, balanced across education and occupations), communication preferences (language complexity, detail level, style), expertise levels (beginner–expert across finance, health, education, technology), behavioral patterns (patience, engagement, multitasking), and emotional profiles (stress, confidence, response speed). This design ensures scalability, reproducibility, diversity without recruitment bias, and ethical efficiency (privacy, consent, fatigue).

### v2

@@ Abstract
Current conversational AI systems often provide generic, one-size-fits-all interactions that overlook individual user characteristics and lack adaptive dialogue management. To address this gap, we introduce HumAIne-chatbot, an AI-driven conversational agent that personalizes responses through a novel user profiling framework. The system is pre-trained on a diverse set of GPT-generated virtual personas to establish a broad prior over user types. During live interactions, an online reinforcement learning agent refines per-user models by combining implicit signals (e.g. typing speed, sentiment, engagement duration) with explicit feedback (e.g., likes and dislikes). This profile dynamically informs the chatbot dialogue policy, enabling real-time adaptation of both content and style. To evaluate the system, we performed controlled experiments with 50 synthetic personas in multiple conversation domains. The results showed consistent improvements in user satisfaction, personalization accuracy, and task achievement when personalization features were enabled. Statistical analysis confirmed significant differences between personalized and nonpersonalized conditions, with large effect sizes across key metrics. These findings highlight the effectiveness of AI-driven user profiling and provide a strong foundation for future real-world validation.
@@ III-C4 AI-Driven User Profiler with Reference Data Integration
Phase I: Virtual Persona Pre‑training
We leverage the Anthology of GPT‑generated backstories to simulate a diverse set of virtual personas. Each backstory encodes rich demographic and behavioral traits (age, occupation, language style), conditioning an LLM to "play" a consistent user archetype [32]. By engaging these personas in multi‑turn dialogues, we collect a large synthetic corpus with "ground‑truth" profile labels. The profiler is first trained in a supervised manner on this corpus, learning to infer latent user attributes and preferences from conversational cues.
This approach aligns with the VirtualXAI framework, which uses GPT‑generated personas for qualitative assessment of explainability methods [33].
@@
We adopt Proximal Policy Optimization (PPO) [36] due to its stability and sample efficiency in high‑dimensional, continuous action spaces. PPO updates policy parameters via clipped objective functions, ensuring gradual behavioral shifts—essential for maintaining user trust during personalization. The actor‑critic architecture computes advantage estimates using a value network baseline, reducing variance in gradient estimates.
@@
Our two‑stage framework builds on recent work combining persona simulation and RL. Wang et al. [35] demonstrate efficacy of a supervised‑to‑RL pipeline for realistic user simulators, while Li et al. [34] show that policy gradient methods improve dialogue coherence and engagement.
@@ IV Evaluation Strategy
To validate HumAIne‐chatbot, we adopted a virtual persona–based evaluation approach inspired by VirtualXAI framework [37] that has used the approached proposed by [13]. Instead of human recruitment, we generated 50 diverse virtual personas with systematically varied demographics (18–65+, balanced across education and occupations), communication preferences (language complexity, detail level, style), expertise levels (beginner–expert across finance, health, education, technology), behavioral patterns (patience, engagement, multitasking), and emotional profiles (stress, confidence, response speed). This design ensures scalability, reproducibility, diversity without recruitment bias, and ethical efficiency (privacy, consent, fatigue).